Part C: Price Elasticity of Demand 10. Given the following demand schedule, calculate the price elasticity of demand for a price change from $40 to $35. Use the midpoint formula and show all work for full credit. (2 points) Price (S) 45 40 35 30 25 20 15 10 Quantity Demanded 15 30 45 60 75 90 105 120 135 11. Using the schedule above, calculate the elasticity of demand when price changes from $25 to $20. Again, show all of your work. Why are the calculated clasticities different? What factors may explain these differences. (2 points) 12. Farmer Bob sells organic tomatoes at the Haile Village Farmers Market every week TAfter talking with some of his customers, he believes that by lowering the price of tomatoes, he can increase total revenue. What is Farmer Bob assuming about the price elasticity of demand for organic tomatoes? What features of Farmer Bobs product might inform this assumption? (2 points) 13, Given the following own-price elasticities, what do you know about the good? Be sure to interpret the elasticities in your explanation. (4 points) Own Price ElasticityTA 0.58 Good 1 Asparagus Cantaloupe Milk Bottled Water 0.1
Answer #1

10) Here, P1 = 40, Q1 = 15,   P2 = 35,   Q2 = 30

PED = (Q2 - Q1) / (P2 - P1) * (P1 + P2) / (Q1 + Q2)

         = (30 - 15) / (35 - 40) * (40 + 35) / (15 + 30)

         = (15 / -5) * (75 / 45)

         = - 5

The absolute value of PED is 5.

11. Here, P1 = 25, Q1 = 60,   P2 = 20,   Q2 = 75

PED = (Q2 - Q1) / (P2 - P1) * (P1 + P2) / (Q1 + Q2)

         = (75 - 60) / (20 - 25) * (25 + 20) / (60 + 75)

         = (15 / -5) * (45 / 135)

         = - 1

The absolute value of PED is 1.

Elasticity measures the relative response of quantity to changes in price. Here, these relative response of quantity to changes in price are different.

Following are the factors on which elasticity of demand depends:

  • Number of substitutes available
  • Proportion of income spent on a good
  • Number of uses of a good
  • Complementary relationship between goods.